Two thirds of the students in a class are boys and the rest are girls. It is known that the probability of a girl getting first class is 0.25 and that of a boy is getting a first class is 0.28. Find the probability that a student chosen at random will get first class marks in the subject

Grade:12

1 Answers

Piyush Kumar Behera
7 years ago
given that in a  class 2/3 are boys and 1/3 are girls
P(B)=2/3
P(G)=1/3
A student chosen at random may be boy or a girl
the required probability=P(B)x0.28+P(G)x0.25
                                 =0.27